# Faster Model Predictive Control via Self-Supervised Initialization Learning

### v2

@@ IV-C Experiment Results
  • MPC with warm-start policy trained by either offline learning or a combination of offline and online fine-tuning significantly outperforms those with all-zero or previous solution-derived guesses, crucial for navigating sharp turns.
  • The offline training algorithm significantly warm-starts the online training algorithm and make it take much less training time to get a satisfying performance.
  • The warm-start policy trained with our two-phase learning framework reduces MPC optimization time and $xte$ on both training and zero-shot tracks, showing the capability of our online fine-tuning algorithm in addressing the suboptimality problem.
  • The two-phase training algorithm outperforms in zero-shot scenarios, demonstrating superior generalizability.
V LIMITATIONS AND FUTURE WORKS
While our proposed two-phase learning framework shows promising results in expediting optimization processes and enhancing control performance for robot control tasks, there are still limitations and avenues for future research that merit consideration. Future work could expand into more complex domains, such as robot arm manipulation or quadcopter control, to explore the algorithm’s efficacy in higher-dimensional spaces. Additionally, future research could also involve utilizing our proposed algorithm to expedite other iterative control optimization techniques like MPPI. Although our algorithm primarily interfaces with MPC, it is versatile enough to serve as a warm start for other iterative control optimization methods like Model Predictive Path Integral (MPPI) control [32]. MPPI typically employs stochastic optimization, initializing with a guess of the control sequence distribution, sampling sequences, and iteratively refining the distribution to minimize control cost. Our framework can be trained to predict the trajectory Gaussian distribution mean, enhancing MPPI’s performance by offering a superior initial distribution. This reduces the iteration count needed for optimization, aligning with early stopping criteria. While MPC suits systems with complex dynamics and constraints, we choose it as our focus due to its representativeness in trajectory optimization methods.
VI CONCLUSIONS
In this paper, we introduce a novel approach to accelerate MPC optimization by learning a warm-start policy. Our two-phase framework combines offline BC and online fine-tuning to provide better initial guesses for the MPC solver. Experiments on training and zero-shot tracks demonstrate the effectiveness of our approach in reducing optimization time and improving path tracking accuracy. This integration of learning with MPC enhances the efficiency and applicability of trajectory optimization in dynamic systems.

### v3

@@ V-C Findings
  • Our proposed framework, trained offline or with online fine-tuning, outperform other non-learning-based baselines.
  • The two-phase training algorithm generalizes better on zero-shot tracks, achieving superior zero-shot performance.
  • Offline training effectively accelerates online fine-tuning, reducing training time to achieve strong performance.
VI LIMITATIONS AND FUTURE WORKS
While our proposed two-phase learning framework shows promising results in expediting optimization processes and enhancing control performance for robot control tasks, it also has several limitations for future research that merit consideration. First, our experiments assume access to perfect state information. Second, the current evaluation is limited to vehicle control. Future work could expand into higher-DOF robotic platforms such as manipulators or aerial robots.
VII CONCLUSIONS
In this paper, we introduce a novel approach to accelerate MPC optimization by learning a warm-start policy. Our two-phase framework combines offline BC and online fine-tuning to provide better initial guesses for the MPC solver. Experiments on training and zero-shot tracks demonstrate the effectiveness of our approach in reducing optimization time without degrading MPC’s performance. This integration of learning with MPC enhances the efficiency and applicability of trajectory optimization in dynamic systems.